OpenStudy (anonymous):

(10k ÷ 2) + 4 ≤ 14 8k + 3 = 20 k - 13 > -10 9k + 15 ≤ 31

OpenStudy (anonymous):

The answer to each question is like that. Question 1 : \[\frac{ 10k }{ 2 } + 4 \le 14\] \[5k \le 14 - 4\] \[5k \le 10\] \[k \le2\] Question 2 : \[8k = 20 - 3\] \[8k = 17 \] \[k = \frac{ 17 }{ 8 }\] Question 3 : \[k > -10 + 13\] \[k > 3\] Question 4: \[9k \le 31 -15\] \[9k \le 16\] \[k \le \frac{ 16 }{ 9 }\]